# Can World Simulators Reason? Gen-ViRe: A Generative Visual Reasoning Benchmark

### v1

@@ 1 Introduction
Recent breakthroughs in Chain-of-Thought (CoT) prompting have demonstrated that large-scale language models (LLMs) can exhibit emergent reasoning capabilities across diverse symbolic domains, from mathematical problem-solving to code generation [cot, gpt-o1, deepseek-r1]. These advances represent a significant milestone in AI’s ability to perform complex reasoning. However, CoT operates exclusively in the symbolic realm—reasoning over language tokens about abstract concepts. While this enables powerful logical inference, it fundamentally cannot capture the continuous, visual, and spatial nature of the real world. This limitation is evident in tasks ranging from abstract spatial navigation to complex physical interaction.
@@
This gap—the need for dynamic, generative reasoning—is also precisely where conventional computer vision paradigms fall short. Conventional vision models function as specialized, passive perceivers—excelling at singular discriminative tasks like object detection or segmentation, yet lacking the unified generative reasoning capabilities needed to simulate dynamic processes. Recognizing this gap, recent pioneering works have proposed a revolutionary paradigm shift: video models as zero-shot learners and world simulators [sora, genie, veo, cof]. As world simulators, these models are transitioning from merely rendering pixels to building an implicit spatiotemporal and physical engine by training on massive video data. This emergent “simulator” then enables them to act as zero-shot learners; they are no longer confined to the specific compositions seen during training but can, much like humans, leverage a fundamental understanding of the world to “zero-shot” imagine, reason about, and create entirely new, logical scenarios. At the core of this new paradigm lies the Chain-of-Frames (CoF) mechanism [cof], where reasoning is actualized through frame-by-frame video generation. Unlike CoT’s discrete symbolic transitions (text $\rightarrow$ text), CoF materializes reasoning as a goal-directed, continuous visual state evolution (frame $\rightarrow$ frame). When a model solves a visual Sudoku puzzle, navigates a maze, or plans multi-step tool use through CoF, each generated frame represents an incremental reasoning step that is both physically grounded and temporally coherent. The generative process itself becomes an act of thinking, not just describing the symbolic coordinates for the maze but generating the dynamic trajectory of the red square moving. It transforms models from passive descriptors into executable world simulators.
@@
The potential of this paradigm is evidenced by rapid progress in video generation technology. Large-scale models—from commercial systems like Sora [sora] and Kling [kling] to open models like CogVideo [cogvideo] and HunyuanVideo [hunyuanvideo]—have demonstrated impressive emergent capabilities in understanding physical interactions, temporal causality, and spatial relationships. Qualitative demonstrations from pioneering works are compelling: models generating coherent multi-step processes, respecting object permanence, and exhibiting intuitive physics. Yet despite this progress, a fundamental question remains unanswered: how well do these models actually reason? Qualitative showcases, while impressive, provide no systematic measure of reasoning depth. More critically, existing video generation benchmarks, which focus primarily on fidelity, alignment, or object-level consistency, do not assess CoF reasoning and therefore cannot measure a model’s core cognitive abilities in multi-step planning and algorithmic logic, or abstract pattern extrapolation [huang2024vbench, liu2024evalcrafter, liu2023fetv, fan2024aigcbench, han2025videobench, sun2025t2vcompbench, wang2024longvideobench, wu2024vq2, tu2021videval, li2023mvbench, chen2023videodreambench, kumar2023tempo, doersch2022tapvid, zhang2024mmevideo]. Without rigorous quantitative evaluation, we cannot distinguish genuine reasoning from sophisticated pattern matching, nor can we diagnose where models fail—perception, physics understanding, or planning—to guide systematic improvement.
@@
To address this critical evaluation gap, we propose Gen-ViRe (Generative VisualReasoning Benchmark), the first comprehensive framework purpose-built to systematically assess CoF reasoning by video generation models. Gen-ViRe is grounded in dual foundations: cognitive science principles defining core pillars of human reasoning [spelke2007coreknowledge, kahneman2011thinking, gentner1983structuremapping, johnsonlaird1983mentalmodels, raven2000ravenreview, spelke2012origins, newell1972humanproblemsolving, byrne2007rationalimagination] , and application-driven requirements from emerging domains like embodied AI and autonomous systems [Feng2025EmbodiedAI, Zhang2025EmbodiedVSR]. We decompose Generative VisualReasoning into six fundamental dimensions— Perceptual, Analogical, Algorithmic & Logical, Spatio-Temporal & Dynamic, Procedural & Planning, and Abstract Reasoning—spanning the complete spectrum from foundational perception to high-order planning [Liang2025AIReasoning, NeuroSymbolicSurvey2023, BoberIrizar2024NeuralAbstraction]. Our data curation employs a multi-source strategy as shown in Figure 4: curating from web resources and academic papers, integrating existing datasets (e.g., ARC-AGI for abstract reasoning, KiVA for analogical tasks), and generating novel tasks using T2I models where suitable data is absent. We implement rigorous prompt design through minimal prompting principles—deliberately providing only high-level goals to assess autonomous reasoning rather than instruction-following—validated through iterative peer review. Our evaluation methodology features a hybrid VLM-assisted approach: Image VLMs judge final outputs for static reasoning tasks, while Video VLMs assess the complete generation process for dynamic tasks, with each judge provisioned with detailed, sub-category-specific criteria refined through multi-round human validation. Through extensive experiments evaluating 7 state-of-the-art video generation models across 72 distinct reasoning prompts (with each model generating 5 instances per prompt, totaling 360 videos per model and over 2,500 overall), Gen-ViRe delivers the first systematic quantification of CoF reasoning capabilities, revealing critical gaps between qualitative potential and quantitative performance, and establishing scientific baselines to guide the development of genuine world simulators.
@@ 2 Related Work
Chain-of-Thought (CoT) prompting has revolutionized how large language models approach complex reasoning tasks by explicitly generating intermediate reasoning steps [cot]. GPT-o1 [gpt-o1] demonstrated that LLMs can leverage CoT for test-time scaling—trading computation for reasoning depth—while DeepSeek-R1 [deepseek-r1] advanced this through Reinforcement Learning with Verifiable Rewards (RLVR), democratizing sophisticated reasoning via open-source release. Recent work has extended CoT to multimodal domains: GPT-4o [gpt-4o] pioneered vision-language joint reasoning, and Bagel [bagel] leverages vision-text interleaved CoT to enhance visual generation. However, CoT remains fundamentally symbolic and discrete—operating in the space of language tokens rather than continuous visual states. While effective for formal logic, CoT is inherently discriminative and passive: it cannot dynamically simulate the physical evolution of the visual world. A CoT chain can describe "a ball rolls down a slope," but cannot generate the actual frame-by-frame trajectory governed by gravity and collision. This limitation necessitates a new paradigm grounded in continuous, generative visual simulation.
@@
The emerging Chain-of-Frames (CoF) paradigm represents a revolutionary shift in how AI systems perform reasoning [cof]. Unlike CoT, which operates on discrete symbolic transitions (text → text), CoF materializes reasoning as continuous visual state evolution (frame → frame). In this paradigm, reasoning is not merely described—it is executed and visualized through frame-by-frame video generation. When a model generates a sequence showing how to solve a visual puzzle, navigate a maze, or manipulate objects to achieve a goal, each generated frame represents an incremental reasoning step that is physically grounded and temporally coherent. Recent pioneering works have begun to explore this new frontier. The Genie series [genie] demonstrates how interactive world generation can enable embodied reasoning, allowing agents to learn causal relationships through simulated physical interactions. Veo-3 pushes the boundaries of long-horizon reasoning, generating extended video sequences that maintain spatial-temporal consistency across complex dynamic scenes [veo]. These works provide compelling qualitative evidence that video models have the potential as world simulators, capable of understanding physics, causality, and spatial relationships [cof].
@@
Recent years have witnessed remarkable progress in video generation, predominantly driven by diffusion-based architectures. Models such as Sora [sora], Kling [kling], and SeeDance [seedance] have achieved breakthroughs in visual quality, generation length, and resolution through large-scale diffusion transformers. While these commercial models remain closed-source, the research community has also contributed open-sourced models like CogVideo [cogvideo] and HunyuanVideo [hunyuanvideo] and WanVideo [wanvideo]. These models demonstrate impressive implicit learning of physical principles purely from observing massive video datasets. Crucially, these models are evolving beyond mere visual generators—they are becoming potential world simulators. For instance, when prompted to generate a video of “a basketball bouncing downstairs”, Veo-3 [veo] can produce sequences that respect gravity, conserve momentum, and exhibit realistic deformation upon impact [cof].
@@ 4.1 Data Collection
Web and Academic Sources. The foundation of our data collection comes from public web resources and academic publications. We use targeted keywords to collect candidate images from search engines like Google. Concurrently, we extract high-quality figures, illustrations, and qualitative examples from relevant academic papers [gu2024controlvla] [wiedemer2024testtime]. Notably, for the Perceptual Reasoning category, we also sourced numerous children’s intelligence tests from the web (e.g., color, shape, and quantity matching puzzles). All collected materials underwent rigorous manual screening and secondary editing (e.g., cropping, annotation, content modification)—such as the modifications made to the aforementioned Spatial Obstacle image—to ensure they perfectly align with the logical and visual constraints of our benchmark tasks.
Integration of Existing Datasets. To evaluate model capabilities in specific domains, we extracted or adapted tasks from several public datasets. This includes GUI navigation datasets [screenspotpro] for Planning Reasoning, Geometry datasets [geolaux] for Algorithmic & Logical Reasoning, and tasks borrowed from KiVA [kiva] for Analogical Reasoning. Critically, to test high-order abstract reasoning, we also incorporated challenging tasks from the ARC-AGI benchmark [ARC-AGI], which is widely considered a gold standard for evaluating fluid intelligence.
Generative Data Creation. For many tasks in our benchmark—especially in the Planning Reasoning category (e.g., Tool Selection and Use)—no large-scale, logically-consistent datasets readily exist. Inspired by the qualitative examples in the pioneering work on Chain-of-Frames [cof], we defined the generative rules and underlying logic for these advanced reasoning tasks and leveraged advanced Text-to-Image models to create entirely new visual puzzles. This approach allows us to systematically control task difficulty, compositional complexity, and generalization requirements, which is unachievable by passively collecting existing data.
@@ 5 Evaluation Methodology
VLM-Assisted Evaluation Methods.   Our evaluation methodology employs powerful Vision Language Models (VLMs) as automated judges. We utilize Gemini 2.5 Pro [Comanici2025Gemini] as our unified VLM judge, leveraging its respective modality capabilities based on the task’s requirements. For tasks where the evaluation is contingent upon the final visual output (such as analogical or geometric reasoning), we employ Gemini 2.5 Pro as an Image VLM judge. For more complex dynamic tasks (such as planning and spatial-temporal reasoning) that require assessing the entire generated process, we utilize Gemini 2.5 Pro as a Video VLM judge. Crucially, for every task, the designated VLM judge (Gemini 2.5 Pro) is provisioned with a detailed, sub-category-specific set of criteria. The judge then decomposes and assesses the model’s output against each criterion, providing an independent score for its decision. This criteria-centric approach ensures a consistent and rigorous evaluation across the entire benchmark.
@@ 6 Experiments
Kling-v1 [kling] 360 0.198 0.071 0.057 0.117 0.140 0.443 0.359
Seedance-1.0-Lite [seedance] 360 0.279 0.087 0.256 0.083 0.146 0.572 0.532
Seedance-1.0-Pro [seedance] 360 0.301 0.154 0.164 0.083 0.171 0.609 0.621
Wan-2.5 [wanvideo] 360 0.490 0.412 0.411 0.500 0.378 0.702 0.536
Veo-3.1 [veo] 360 0.486 0.440 0.451 0.367 0.386 0.722 0.550
Hailuo-2.3 [hailuo] 360 0.493 0.494 0.355 0.383 0.425 0.778 0.524
Sora-2 [sora] 360 0.560 0.604 0.472 0.483 0.496 0.768 0.537
@@ 6.1 Experimental Setup
Evaluation Metrics. The VLM-assisted evaluation process described in Section 5 yields an output score for each task. To aggregate these individual scores, we follow the aggregation strategy of MEGA-BENCH [chen2024mega]. The score for each individual task is first normalized to a consistent [0, 1] range, where 1.0 signifies perfect adherence to all standards. Subsequently, to report comprehensive performance, we compute the macro-mean score of all normalized scores.
@@ 7 Conclusion
[…]
Supplementary Material
8 Rationale
Having the supplementary compiled together with the main paper means that:
  • The supplementary can back-reference sections of the main paper, for example, we can refer to LABEL:sec:intro;
  • The main paper can forward reference sub-sections within the supplementary explicitly (e.g. referring to a particular experiment);
  • When submitted to arXiv, the supplementary will already included at the end of the paper.
To split the supplementary pages from the main paper, you can use Preview (on macOS), Adobe Acrobat (on all OSs), as well as command line tools.
8.1 Qualitative Analysis and Case Studies
Analogical Reasoning (Rotation and Color)
Algorithmic & Logical Reasoning (Geometry)
Figure 11: Showcase of Sora-2’s in geometry task. Both Sora-2 and Veo-3.1 failed to identify the existing point D in the static image. This perceptual error caused them to ignore the instruction ’connect point C to point D’ and instead hallucinate a new point D, connecting C to this incorrect location.
I2V Prompt. A static view. Draw a line connecting point C to point D, and end the video once the line is fully drawn.
Algorithmic & Logical Reasoning (Sudoku)
Figure 12: Showcase of Sora-2’s Sudoku task. In the Sudoku task, Sora-2 exhibits an emergent, human-like thinking process. The model uses a question mark (?) as a placeholder for the unknown value in the third row. This suggests it can hold an internal state of the problem ("this cell is unsolved"). Following the placeholder, the model generates frames that simulate the "moving" of numbers (2) into their correct, logically-deduced positions to complete the row.
Takeaway. Sora-2’s Sudoku solving process (using "?" placeholders and "moving" numbers) indicates it is acquiring a genuine algorithmic capability, as it is simulating the problem-solving process of "following Sudoku rules" rather than just pattern-matching the final answer.
Spatio-Temporal Reasoning (Spatial Obstacle)
Takeaway.Sora-2’s performance reveals a critical divergence: abstract logical reasoning (e.g., following algorithmic rules) and physical reality simulation (e.g., adhering to physical plausibility) are two distinct capabilities. The model’s excellence in the former does not equate to mastery in the latter.

### v2

@@ 1 Introduction
Recent breakthroughs in Chain-of-Thought (CoT) prompting have demonstrated that large-scale language models (LLMs) can exhibit emergent reasoning capabilities across diverse symbolic domains, from mathematical problem-solving to code generation [43, 23, 17]. These advances represent a significant milestone in AI’s ability to perform complex reasoning. However, CoT operates exclusively in the symbolic realm—reasoning over language tokens about abstract concepts. While this enables powerful logical inference, it fundamentally cannot capture the continuous, visual, and spatial nature of the real world. This limitation is evident in tasks ranging from abstract spatial navigation to complex physical interaction.
@@
This gap—the need for dynamic, generative reasoning—is also precisely where conventional computer vision paradigms fall short. Conventional vision models function as specialized, passive perceivers—excelling at singular discriminative tasks like object detection or segmentation, yet lacking the unified generative reasoning capabilities needed to simulate dynamic processes. Recognizing this gap, recent pioneering works have proposed a revolutionary paradigm shift: video models as zero-shot learners and world simulators [2, 3, 15, 46]. As world simulators, these models are transitioning from merely rendering pixels to building an implicit spatiotemporal and physical engine by training on massive video data. This emergent “simulator” then enables them to act as zero-shot learners; they are no longer confined to the specific compositions seen during training but can, much like humans, leverage a fundamental understanding of the world to “zero-shot” imagine, reason about, and create entirely new, logical scenarios. At the core of this new paradigm lies the Chain-of-Frames (CoF) mechanism [46], where reasoning is actualized through frame-by-frame video generation. Unlike CoT’s discrete symbolic transitions (text $\rightarrow$ text), CoF materializes reasoning as a goal-directed, continuous visual state evolution (frame $\rightarrow$ frame). When a model solves a visual Sudoku puzzle, navigates a maze, or plans multi-step tool use through CoF, each generated frame represents an incremental reasoning step that is both physically grounded and temporally coherent. The generative process itself becomes an act of thinking, not just describing the symbolic coordinates for the maze but generating the dynamic trajectory of the red square moving. It transforms models from passive descriptors into executable world simulators.
@@
The potential of this paradigm is evidenced by rapid progress in video generation technology. Large-scale models—from commercial systems like Sora [2] and Kling [28] to open models like CogVideo [19] and HunyuanVideo [27]—have demonstrated impressive emergent capabilities in understanding physical interactions, temporal causality, and spatial relationships. Qualitative demonstrations from pioneering works are compelling: models generating coherent multi-step processes, respecting object permanence, and exhibiting intuitive physics. Yet despite this progress, a fundamental question remains unanswered: how well do these models actually reason? Qualitative showcases, while impressive, provide no systematic measure of reasoning depth. More critically, existing video generation benchmarks, which focus primarily on fidelity, alignment, or object-level consistency, do not assess CoF reasoning and therefore cannot measure a model’s core cognitive abilities in multi-step planning and algorithmic logic, or abstract pattern extrapolation [20, 33, 32, 11, 18, 39, 42, 47, 40, 30, 6, 29, 10, 50]. Without rigorous quantitative evaluation, we cannot distinguish genuine reasoning from sophisticated pattern matching, nor can we diagnose where models fail—perception, physics understanding, or planning—to guide systematic improvement.
@@
To address this critical evaluation gap, we propose Gen-ViRe (Generative VisualReasoning Benchmark), the first comprehensive framework purpose-built to systematically assess CoF reasoning by video generation models. Gen-ViRe is grounded in dual foundations: cognitive science principles defining core pillars of human reasoning [37, 26, 14, 25, 36, 38, 35, 4] , and application-driven requirements from emerging domains like embodied AI and autonomous systems [12, 49]. We decompose Generative VisualReasoning into six fundamental dimensions— Perceptual, Analogical, Algorithmic & Logical, Spatio-Temporal & Dynamic, Procedural & Planning, and Abstract Reasoning—spanning the complete spectrum from foundational perception to high-order planning [31, 48, 1]. Our data curation employs a multi-source strategy as shown in Figure 4: curating from web resources and academic papers, integrating existing datasets (e.g., ARC-AGI for abstract reasoning, KiVA for analogical tasks), and generating novel tasks using T2I models where suitable data is absent. We implement rigorous prompt design through minimal prompting principles—deliberately providing only high-level goals to assess autonomous reasoning rather than instruction-following—validated through iterative peer review. Our evaluation methodology features a hybrid VLM-assisted approach: Image VLMs judge final outputs for static reasoning tasks, while Video VLMs assess the complete generation process for dynamic tasks, with each judge provisioned with detailed, sub-category-specific criteria refined through multi-round human validation. Through extensive experiments evaluating 7 state-of-the-art video generation models across 72 distinct reasoning prompts (with each model generating 5 instances per prompt, totaling 360 videos per model and over 2,500 overall), Gen-ViRe delivers the first systematic quantification of CoF reasoning capabilities, revealing critical gaps between qualitative potential and quantitative performance, and establishing scientific baselines to guide the development of genuine world simulators.
@@ 2 Related Work
Chain-of-Thought (CoT) prompting has revolutionized how large language models approach complex reasoning tasks by explicitly generating intermediate reasoning steps [43]. GPT-o1 [23] demonstrated that LLMs can leverage CoT for test-time scaling—trading computation for reasoning depth—while DeepSeek-R1 [17] advanced this through Reinforcement Learning with Verifiable Rewards (RLVR), democratizing sophisticated reasoning via open-source release. Recent work has extended CoT to multimodal domains: GPT-4o [21] pioneered vision-language joint reasoning, and Bagel [9] leverages vision-text interleaved CoT to enhance visual generation. However, CoT remains fundamentally symbolic and discrete—operating in the space of language tokens rather than continuous visual states. While effective for formal logic, CoT is inherently discriminative and passive: it cannot dynamically simulate the physical evolution of the visual world. A CoT chain can describe "a ball rolls down a slope," but cannot generate the actual frame-by-frame trajectory governed by gravity and collision. This limitation necessitates a new paradigm grounded in continuous, generative visual simulation.
@@
The emerging Chain-of-Frames (CoF) paradigm represents a revolutionary shift in how AI systems perform reasoning [46]. Unlike CoT, which operates on discrete symbolic transitions (text → text), CoF materializes reasoning as continuous visual state evolution (frame → frame). In this paradigm, reasoning is not merely described—it is executed and visualized through frame-by-frame video generation. When a model generates a sequence showing how to solve a visual puzzle, navigate a maze, or manipulate objects to achieve a goal, each generated frame represents an incremental reasoning step that is physically grounded and temporally coherent. Recent pioneering works have begun to explore this new frontier. The Genie series [3] demonstrates how interactive world generation can enable embodied reasoning, allowing agents to learn causal relationships through simulated physical interactions. Veo-3 pushes the boundaries of long-horizon reasoning, generating extended video sequences that maintain spatial-temporal consistency across complex dynamic scenes [15]. These works provide compelling qualitative evidence that video models have the potential as world simulators, capable of understanding physics, causality, and spatial relationships [46].
@@
Recent years have witnessed remarkable progress in video generation, predominantly driven by diffusion-based architectures. Models such as Sora [2], Kling [28], and SeeDance [13] have achieved breakthroughs in visual quality, generation length, and resolution through large-scale diffusion transformers. While these commercial models remain closed-source, the research community has also contributed open-sourced models like CogVideo [19] and HunyuanVideo [27] and WanVideo [41]. These models demonstrate impressive implicit learning of physical principles purely from observing massive video datasets. Crucially, these models are evolving beyond mere visual generators—they are becoming potential world simulators. For instance, when prompted to generate a video of “a basketball bouncing downstairs”, Veo-3 [15] can produce sequences that respect gravity, conserve momentum, and exhibit realistic deformation upon impact [46].
@@ 4.1 Data Collection
Web and Academic Sources. The foundation of our data collection comes from public web resources and academic publications. We use targeted keywords to collect candidate images from search engines like Google. Concurrently, we extract high-quality figures, illustrations, and qualitative examples from relevant academic papers [16] [45]. Notably, for the Perceptual Reasoning category, we also sourced numerous children’s intelligence tests from the web (e.g., color, shape, and quantity matching puzzles). All collected materials underwent rigorous manual screening and secondary editing (e.g., cropping, annotation, content modification)—such as the modifications made to the aforementioned Spatial Obstacle image—to ensure they perfectly align with the logical and visual constraints of our benchmark tasks.
Integration of Existing Datasets. To evaluate model capabilities in specific domains, we extracted or adapted tasks from several public datasets. This includes GUI navigation datasets [22] for Planning Reasoning, Geometry datasets [44] for Algorithmic & Logical Reasoning, and tasks borrowed from KiVA [24] for Analogical Reasoning. Critically, to test high-order abstract reasoning, we also incorporated challenging tasks from the ARC-AGI benchmark [7], which is widely considered a gold standard for evaluating fluid intelligence.
Generative Data Creation. For many tasks in our benchmark—especially in the Planning Reasoning category (e.g., Tool Selection and Use)—no large-scale, logically-consistent datasets readily exist. Inspired by the qualitative examples in the pioneering work on Chain-of-Frames [46], we defined the generative rules and underlying logic for these advanced reasoning tasks and leveraged advanced Text-to-Image models to create entirely new visual puzzles. This approach allows us to systematically control task difficulty, compositional complexity, and generalization requirements, which is unachievable by passively collecting existing data.
@@ 5 Evaluation Methodology
VLM-Assisted Evaluation Methods.   Our evaluation methodology employs powerful Vision Language Models (VLMs) as automated judges. We utilize Gemini 2.5 Pro [8] as our unified VLM judge, leveraging its respective modality capabilities based on the task’s requirements. For tasks where the evaluation is contingent upon the final visual output (such as analogical or geometric reasoning), we employ Gemini 2.5 Pro as an Image VLM judge. For more complex dynamic tasks (such as planning and spatial-temporal reasoning) that require assessing the entire generated process, we utilize Gemini 2.5 Pro as a Video VLM judge. Crucially, for every task, the designated VLM judge (Gemini 2.5 Pro) is provisioned with a detailed, sub-category-specific set of criteria. The judge then decomposes and assesses the model’s output against each criterion, providing an independent score for its decision. This criteria-centric approach ensures a consistent and rigorous evaluation across the entire benchmark.
@@ 6 Experiments
Kling-v1 [28] 360 0.198 0.071 0.057 0.117 0.140 0.443 0.359
Seedance-1.0-Lite [13] 360 0.279 0.087 0.256 0.083 0.146 0.572 0.532
Seedance-1.0-Pro [13] 360 0.301 0.154 0.164 0.083 0.171 0.609 0.621
Wan-2.5 [41] 360 0.490 0.412 0.411 0.500 0.378 0.702 0.536
Veo-3.1 [15] 360 0.486 0.440 0.451 0.367 0.386 0.722 0.550
Hailuo-2.3 [34] 360 0.493 0.494 0.355 0.383 0.425 0.778 0.524
Sora-2 [2] 360 0.560 0.604 0.472 0.483 0.496 0.768 0.537
@@ 6.1 Experimental Setup
Evaluation Metrics. The VLM-assisted evaluation process described in Section 5 yields an output score for each task. To aggregate these individual scores, we follow the aggregation strategy of MEGA-BENCH [5]. The score for each individual task is first normalized to a consistent [0, 1] range, where 1.0 signifies perfect adherence to all standards. Subsequently, to report comprehensive performance, we compute the macro-mean score of all normalized scores.